### v1

# Efficient Digital Quadratic Unconstrained Binary Optimization Solvers for SAT Problems

Abstract
Boolean satisfiability (SAT) is a propositional logic problem of determining whether an assignment of variables satisfies a Boolean formula. Many combinatorial optimization problems can be formulated in Boolean SAT logic — either as $k$-SAT decision problems or Max $k$-SAT optimization problems, with conflict-driven (CDCL) solvers being the most prominent. Despite their ability to handle large instances, CDCL-based solvers have fundamental scalability limitations. In light of this, we propose recently-developed quadratic unconstrained binary optimization (QUBO) solvers as an alternative platform for $3$-SAT problems. To utilize them, we implement a $2$-step [$3$-SAT]-[Max $2$-SAT]-[QUBO] conversion procedure and present a rigorous proof to explicitly calculate the number of both satisfied and violated clauses of the original $3$-SAT instance from the transformed Max $2$-SAT formulation. We then demonstrate, through numerical simulations on several benchmark instances, that digital QUBO solvers can achieve state-of-the-art accuracy on $78$-variable $3$-SAT benchmark problems. Our work facilitates the broader use of quantum annealers on noisy intermediate-scale quantum (NISQ) devices, as well as their quantum-inspired digital counterparts, for solving $3$-SAT problems.
preprint: APS/123-QED
I Introduction
The Boolean $k$-SAT problem (for $k\geq 3$) is an NP-complete [1] combinatorial optimization problem with many commercial applications [2, 3, 4, 5, 6, 7], making it an invaluable tool for addressing complex decision-making and problem-solving tasks. SAT problems are solved by determining whether an assignment of Boolean variables, either TRUE or FALSE, satisfies a propositional Boolean formula in conjunctive normal form (CNF), such that the formula evaluates to TRUE. Fundamentally, this is characterized by search for an optimal configuration (e.g. minimum of an objective function) of variables within a vast but finite solution space.
Consequently, most SAT solvers utilize search algorithms in conjunction with decision and branching heuristics [8, 9]. Where the CDCL framework [10], due to its proven efficiency in solving industrial SAT problems [11], has emerged as the backbone of modern solvers. Typically, high-performance CDCL solvers employ a multitude of sophisticated clause minimization [12], distance [13], and history-based [14] heuristics, tailored for crafted benchmarks. Despite accurately solving larger instances, CDCL solvers have inherent scaling constraints [15]. While some learning-focused search techniques exist, attempts to parallelize CDCL solvers often lead to excessive memory usage, which degrades the performance and can render them unsuitable for some inherently large-scale or computationally-hard problems [16, 17].
The ubiquitous nature of SAT problems prompted the search for quantum computing solutions, where recent hardware advancements [18] enabled the embedding of the QUBO form of many discrete optimization problems (e.g. SAT) [19] to quantum annealing devices (see, e.g., [20]), such that the objective becomes minimizing a quadratic polynomial over binary variables. [^1] By virtue of the $3$-SAT problem being NP-complete, any solver that can efficiently solve it can be also used for all problems in the NP class with only polynomial time to verify solutions. Moreover, every SAT problem can be expressed as a 3-SAT problem in CNF [21]. The hope is that, due to their noise resilience (compared to some gate-based models), they can be applied on problems with non-trivial logic structure in the NISQ era. Still, despite continued efforts, demonstrating a clear scaling or accuracy advantage remains challenging.
Regardless, the main underlying principles of adiabatic quantum computing remain valid, thus stimulating the development of digital alternatives; so-called quantum-inspired QUBO solvers. Indeed, they recently emerged as a viable alternative due to several desirable features, such as parameter tuning flexibility and inherent capacity to effectively handle large instances in parallel, see [22]. To that end, progress in ground state convergence [23], energy landscape exploration [24], and multidimensional function minimization [25] bolster the ability of QUBO solvers to successfully tackle discrete optimization problems.
In the absence of full-fledged quantum computers, QUBO solvers can represent a prominent general-purpose optimization platform. The present work is a step in this direction as we propose a $2$-step [$3$-SAT]-[Max $2$-SAT]-[QUBO] conversion procedure for embedding $3$-SAT problems of arbitrary size to available QUBO solvers (see Section II and Appendix A and B for details). By construction, the conversion allows us to retrieve the solution of the original $3$-SAT problem from the QUBO solution. Then, in Section III, we present a rigorous proof to explicitly retrieve the exact number of both satisfied and violated clauses of the original 3-SAT instance from the transformed Max $2$-SAT formulation. Finally, in Section IV, we conduct numerical simulations on several publicly available benchmark instances of varying sizes and clause densities, and show that QUBO solvers can indeed match the accuracy of RC2 [26] — a CDCL-based state-of-the-art Max SAT solver. To the best of our knowledge, this is the first direct performance-oriented comparison of commercial QUBO solvers to RC2222, and fills an important gap in the literature. Our results offer theoretical advancements in the field of discrete optimization and establish the QUBO platform as a high-performance alternative to CDCL for solving $3$-SAT problems on near-term devices.
II Converting 3-SAT problems to QUBO
To use available digital QUBO solvers and take advantage of their inherent properties, such as parameter tuning flexibility and scalability, the logical structure of the original $3$-SAT problem must be converted to the appropriate form, see [19]. Here, we implement a $2$-step [$3$-SAT]-[Max $2$-SAT]-[QUBO] conversion with the so-called $(7,10)$-gadget [27]. For that reason, we herein define “QUBO” as any heuristic that can efficiently solve hard $3$-SAT problems, whose use is enabled by the $(7,10)$-gadget described in [27] and Section II.1 below. Moreover, the presented method is necessary since the $3$-SAT Hamiltonian does not have an Ising quantization, see Appendix A for a detailed proof.
II.1 3-SAT to Max 2-SAT
Consider an arbitrary $3$-SAT problem $\dagger$ with $M$ clauses and $N$ literals expressed in CNF, that is a set of logical AND operations on a string of $M$ disjunctive $3$-literal clauses $\left\{\Omega_{k}\right\}_{k=1}^{M}$ each expressed as logical OR operations on its $N$ literals $\left\{\sigma_{i}\right\}_{i=1}^{N}$. Then we can write $\dagger=\Omega_{1}\land\Omega_{2}\land\dots\land\Omega_{M}$ with the $k^{th}$ clause expressed as:
$\Omega_{k}=\left(\operatorname{sgn}\left(\sigma^{k}_{j_{1}}\right)\sigma^{k}_{j_{1}}\lor\operatorname{sgn}\left(\sigma^{k}_{j_{2}}\right)\sigma^{k}_{j_{2}}\lor\operatorname{sgn}\left(\sigma^{k}_{j_{3}}\right)\sigma^{k}_{j_{3}}\right),$
where $k\in\{1,2,\dots,M\}$, $j_{1},j_{2},j_{3}\in\left\{1,\ldots,N\right\}$, and $\operatorname{sgn}\left(\sigma^{k}_{j_{i}}\right)\sigma^{k}_{j_{i}}=\sigma^{k}_{j_{i}}\text{ or }\neg\sigma^{k}_{j_{i}}$, where equivalently, $\operatorname{sgn}\left(\sigma^{k}_{j_{i}}\right)=\mathtt{TRUE}$ and $\mathtt{FALSE}$, respectively. For the rest of the paper we will, without loss of generality, assume $\dagger$ consists of clauses with exactly $3$ literals.
Without loss of generality, we now illustrate how the $(7,10)$-gadget [27] converts a $3$-SAT problem to Max $2$-SAT with a simple clause $\Omega_{k}=\left(\sigma^{k}_{j_{1}}\lor\sigma^{k}_{j_{2}}\lor\sigma^{k}_{j_{3}}\right)$, where the construction extends naturally to clauses with negations of literals. First, we define exactly one ancillary variable $d^{k}$, and then replace $\Omega_{k}$ by the corresponding set of $10$ Max $2$-SAT clauses:
$\Omega^{\prime}_{k}:=$$\left\{(\sigma^{k}_{j_{1}}),(\sigma^{k}_{j_{2}}),(\sigma^{k}_{j_{3}}),(d^{k}),\right.$
($(\neg\sigma^{k}_{j_{1}}\lor\neg\sigma^{k}_{j_{2}}),(\neg\sigma^{k}_{j_{1}}\lor\neg\sigma^{k}_{j_{3}}),(\neg\sigma^{k}_{j_{2}}\lor\neg\sigma^{k}_{j_{3}}),$
$\left.({\sigma}^{k}_{j_{1}}\lor\neg d^{k}),({\sigma}^{k}_{j_{2}}\lor\neg d^{k}),({\sigma}^{k}_{j_{3}}\lor\neg d^{k})\right\}.$ (1)
After replacing every clause of the original $3$-SAT problem in this manner, the conversion of the of the entire $3$-SAT problem can therefore be expressed compactly as:
[…]
$\bigwedge_{k=1}^{M}\Omega_{k}$ $\mapsto\bigwedge_{k=1}^{M}\Omega^{\prime}_{k}.$
While the index of the converted Max $2$-SAT problem goes to $M$, each $\Omega_{k}^{\prime}$ contains $10$ clauses. Therefore, an arbitrary $3$-SAT problem with $M$ clauses can be converted to a Max $2$-SAT problem with $10M$ clauses and $N+M$ variables, see Table 1.
Evidently, there are $8$ possible choices of $(\sigma^{k}_{j_{1}},\sigma^{k}_{j_{2}},\sigma^{k}_{j_{3}})$, with each being either 00 or $1$. Amongst them, only $(0,0,0)$ will not satisfy $\Omega_{k}$. In particular, given that an assignment $(\sigma^{k}_{j_{1}},\sigma^{k}_{j_{2}},\sigma^{k}_{j_{3}})$ satisfies $\Omega_{k}$, then exactly $7$ clauses of $\Omega^{\prime}_{k}$ will be satisfied. Otherwise, if the initial assignment does not satisfy $\Omega_{k}$, then exactly $6$ clauses of $\Omega^{\prime}_{k}$ will be satisfied. This is summarized in Table 2 in Section III, where the properties of the conversion are studied.
3-SAT Max 2-SAT
Clauses $M$ $10M$
Literals $N$ $N+M$
Table 1: Changes in total number of clauses and literals in the $3$-SAT to Max $2$-SAT conversion (Eq. (II.1)).
II.2 Max 2-SAT to QUBO
The previous Section outlined the conversion of an arbitrary $3$-SAT problem to a Max $2$-SAT problem. Here, we present the second step of the procedure. Consider a Max 2-SAT problem $\ddagger:=\bigwedge_{k=1}^{M^{\prime}}\Omega^{\prime}_{k}$ with $M^{\prime}$ clauses $\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M^{\prime}}$and $N^{\prime}$ literals $\left\{\sigma^{\prime}_{i}\right\}_{i=1}^{N^{\prime}}$, where each clause contains at most two literals. We now define a QUBO problem on $\ddagger$ over the search space $\mathbf{x}\in\left\{0,1\right\}^{N^{\prime}}$, with the goal to minimize the total number of violated clauses. Formally:
We first define a set of variables $\{v_{j}^{k}\}$ over the $M^{\prime}$ clauses. Here, $v_{j}^{k}\in\{-1,0,1\}$, where $v_{j}^{k}=-1(+1)$ if $\sigma^{\prime}_{j}$ is negated (unnegated) in the $k^{th}$ clause, respectively, and $v_{j}^{k}=0$ for all clauses where $\sigma^{\prime}_{j}$ does not appear in the clause. More precisely, for each $\Omega^{\prime}_{k}\in\ddagger$ containing literals with sub-indices $j_{1},j_{2}$, we define variables $\left\{v_{j}^{k}\right\}$ where:
  1. 1.
    If $j=j_{1}$ or $j_{2}$:
    $v_{j}^{k}=2\cdot\operatorname{sgn}\left(\sigma^{\prime k}_{j}\right)-1\Leftrightarrow\begin{cases}1&\operatorname{sgn}\left(\sigma^{\prime k}_{j}\right)=\texttt{TRUE}\\ -1&\operatorname{sgn}\left(\sigma^{\prime k}_{j}\right)=\texttt{FALSE}\end{cases}$
  2. 2.
    $v_{j}^{k}=0$ otherwise.
For each $k\in\{1,\ldots,M^{\prime}\}$, we can isolate the entries $x_{j_{i}}^{k}$ from the input $\mathbf{x}=\left(x_{1},\ldots,x_{j_{1}}^{k},\ldots,x_{j_{2}}^{k},\ldots,x_{N^{\prime}}\right)\in\left\{0,1\right\}^{N^{\prime}}$, and define the following intermediate variable $S_{j_{i}}^{k}$:
$S_{j_{i}}^{k}$ $=\left(\frac{1+v_{j_{i}}^{k}}{2}-v_{j_{i}}^{k}x_{j_{i}}^{k}\right).$
$S^{k}$ $=S_{j_{1}}^{k}\cdot S_{j_{2}}^{k}=\prod_{i=1}^{2}\left(\frac{1+v_{j_{i}}^{k}}{2}-v_{j_{i}}^{k}x_{j_{i}}^{k}\right).$ (2)
One can check that if $\mathbf{x}$ satisfies $\Omega^{\prime}_{k}$, then $S^{k}=0$ and $S^{k}=1$ otherwise. The total number of violated clauses in Max $2$-SAT problem $\ddagger=\bigwedge_{k=1}^{M^{\prime}}\Omega^{\prime}_{k}$, given input $\mathbf{x}$, is:
$\mathcal{V}(\mathbf{x})$ $=\sum_{k=1}^{M^{\prime}}\prod_{i=1}^{2}\left(\frac{1+v_{j_{i}}^{k}}{2}-v_{j_{i}}^{k}x_{j_{i}}^{k}\right)$
$=\sum_{k=1}^{M^{\prime}}\Bigg{(}v_{j_{1}}^{k}v_{j_{2}}^{k}x_{j_{1}}^{k}x_{j_{2}}^{k}-v_{j_{1}}^{k}\frac{1+v_{j_{2}}^{k}}{2}x_{j_{1}}^{k}$
$\quad\quad-v_{j_{2}}^{k}\frac{1+v_{j_{1}}^{k}}{2}x_{j_{2}}^{k}+\frac{1+v_{j_{1}}^{k}}{2}\frac{1+v_{j_{2}}^{k}}{2}\Bigg{)},$
where the optimal solution $\mathbf{x^{*}}$ is thus given by:
$\mathbf{x}^{*}=\mathop{\arg\min}\limits_{\mathbf{x}\in\left\{0,1\right\}^{N^{\prime}}}\mathcal{V}(\mathbf{x}).$ (3)
Equivalently, upon re-scaling $\mathcal{V}(\mathbf{x})$ by a factor of $2$ and dropping the constant term, we obtain the QUBO objective function $q(\mathbf{x})$ corresponding to $\mathcal{V}(\mathbf{x})$, which is to be minimized:
$q(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}\mathbf{Q}\mathbf{x}+\mathbf{b}^{\top}\mathbf{x}+c,$ (4)
where for all $j_{1},j_{2}\in\{1,\ldots,N^{\prime}\}$:
$Q_{j_{1}j_{2}}=2\sum_{k=1}^{M^{\prime}}v_{j_{1}}^{k}v_{j_{2}}^{k}$ (5)
$b_{j_{1}}=-\sum_{k=1}^{M^{\prime}}v_{j_{1}}^{k}\left(v_{j_{2}}^{k}+1\right).$ (6)
For the unassigned $Q_{ij}$ and $b_{i}$ in the above procedure, we set $Q_{ij}=0$, $b_{i}=0$.
III Properties
In this Section we illustrate how information about the original $3$-SAT problem can be extracted from the converted Max $2$-SAT formulation. In particular, we show how to retrieve the exact number of both violated and satisfied clauses of the initial $3$-SAT problem.
We denote the initial $3$-SAT problem by $\dagger:=\bigwedge_{k=1}^{M}\Omega_{k}$ and $\ddagger:=\bigwedge_{k=1}^{M}\Omega^{\prime}_{k}$ the Max $2$-SAT converted from $\dagger$. Recall from Section II.1, for each clause $\Omega_{k}=\left(\sigma^{k}_{j_{1}}\lor\sigma^{k}_{j_{2}}\lor\sigma^{k}_{j_{3}}\right)$ of the original $3$-SAT problem $\dagger$ there are $8$ possible choices of assignments of $\left(\sigma^{k}_{j_{1}},\sigma^{k}_{j_{2}},\sigma^{k}_{j_{3}}\right)$ with $\sigma_{j_{i}}^{k}\in\left\{0,1\right\}$ for $i=1,2,3$. By Theorem 1.1 of [27], $\Omega_{k}$ can be converted to a Max $2$-SAT clause $\Omega_{k}^{\prime}$, where the satisfactory table is outlined in Table 2.
$\left(\sigma^{k}_{j_{1}},\sigma^{k}_{j_{2}},\sigma^{k}_{j_{3}}\right)$ $d^{k}$ $\Omega_{k}^{\prime}$ $\Omega_{k}$
Case 1 $(0,0,0)$ 00 $6$
$1$ $4$
Case 2 $(0,0,1),(0,1,0)$, 00 $7$
$(1,0,0)$ $1$ $6$
Case 3 Otherwise $0,1$ $7$
Table 2: Satisfactory Table of a single clause the $3$-SAT to Max $2$-SAT conversion described by Eq. (II.1). For the column corresponding to $\Omega_{k}^{\prime}$, the label (##\## ✓) denote the number (##\##) of satisfied clause(s). For the column corresponding to $\Omega_{k}$, ✓and ✗denote satisfied and violation respectively.
Hence, we arrive at the following statement, which completes the results in [29]
Lemma III.1.
Given a $3$-SAT clause $\Omega_{k}$ and its corresponding Max $2$-SAT conversion $\Omega_{k}^{\prime}$, there exists a configuration of ancillary variable $d^{k}$ such that: if an assignment satisfies $\Omega_{k}$ then exactly $7$ clauses are satisfied in $\Omega_{k}^{\prime}$, otherwise if an assignment does not satisfy $\Omega_{k}$ then exactly $6$ clauses are satisfied in $\Omega_{k}^{\prime}$.
For the rest of the Section, we show that the exact number of violated and satisfied clauses of the original $3$-SAT problem, for each of the cases outlined in Table 2, can be retrieved from the converted Max $2$-SAT instance once the problem $\ddagger$ is solved. We begin with the special case where the converted Max $2$-SAT problem is solved and all the ancillary variables are identically zero.
Theorem III.2.
Let $\ddagger:=\bigwedge_{k=1}^{M}\Omega^{\prime}_{k}$ denote a Max $2$-SAT problem converted from $3$-SAT problem $\dagger:=\bigwedge_{k=1}^{M}\Omega_{k}$. Suppose $\ddagger$ is solved when the ancillary variables all vanish and the final number of satisfied clauses of $\ddagger$ is known. Then the number of violated and satisfied clauses of $\dagger$ can be retrieved.
Proof.
Suppose $\ddagger$ is solved when the ancillary variables all vanish, i.e. $d^{k}=0$ for all $k=1,\ldots,M$. Let $\mathfrak{S}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses of $\ddagger$, and let $\mathcal{V},\mathcal{S}\in\mathbb{N}_{\geq 0}$ denote the number of violated and satisfied clauses of $\dagger$ respectively. We have the following linear Diophantine equation:
$\mathfrak{S}$ $=6\cdot\mathcal{V}+7\cdot\mathcal{S}.$ (7)
Since $7$ is prime, this equation has integer solutions and can be solved completely by Euclidean Algorithm. Hence, there exists $\hat{\mathcal{V}},\hat{\mathcal{S}}\in\mathbb{Z}$ such that the integer solutions of Eq. (7) take the form:
$\mathcal{V}=\hat{\mathcal{V}}-7\cdot n;$ $\quad\mathcal{S}=\hat{\mathcal{S}}+6\cdot n,\quad\forall n\in\mathbb{Z}.$ (8)
Equivalently and more explicitly, since $\mathcal{V}+\mathcal{S}=M$ by construction, $\mathcal{V},\mathcal{S}\in\mathbb{N}_{\geq 0}$ are the non-negative integer solutions of the following linear system, whose integrality and existence is due to Eq. (8):
$\begin{bmatrix}6&7\\ 1&1\end{bmatrix}\cdot\begin{bmatrix}\mathcal{V}\\ \mathcal{S}\end{bmatrix}=\begin{bmatrix}\mathfrak{S}\\ M\end{bmatrix}.$
By similar argument, in the special case where the converted Max $2$-SAT problem is solved and all the ancillary variables are $1$, the number of violated and sastified clauses can also be retreived from the original $3$-SAT problem.
Corollary III.3.
Let $\ddagger:=\bigwedge_{k=1}^{M}\Omega^{\prime}_{k}$ denote a Max $2$-SAT problem converted from 3-SAT problem $\dagger:=\bigwedge_{k=1}^{M}\Omega_{k}$. Suppose $\ddagger$ is solved when the ancillary variables all equal to $1$ and the final number of satisfied clauses of $\ddagger$ is known. Then the number of violated and satisfied clauses of $\dagger$ can be retrieved.
Proof.
See Appendix C. ∎
Notice that in the general case the clauses of the converted Max $2$-SAT problem $\ddagger$ can be partitioned into two distinctive parts, where the ancillary is either 00 or $1$. [^2] In practice, this can be directly counted from the final solution. We can, therefore, apply Theorem III.2 and Corollary III.3 separately to both parts and recombine the results. In particular, we now show that once $\ddagger$ is solved, the number of violated and satisfied clauses of the original $3$-SAT problem can be retrieved.
Theorem III.4.
Let $\ddagger:=\bigwedge_{k=1}^{M}\Omega^{\prime}_{k}$ denote a Max $2$-SAT problem converted from $3$-SAT problem $\dagger:=\bigwedge_{k=1}^{M}\Omega_{k}$. Suppose $\ddagger$ is solved and the total number of satisfied clauses of $\ddagger$ is counted for both cases when the ancillary variables is 00 or $1$. Then the number of violated and satisfied clauses of $\dagger$, denoted by $\mathcal{V},\mathcal{S}\in\mathbb{N}_{\geq 0}$ respectively, can be retrieved.
Proof.
See Appendix D. ∎
IV Numerical Experiments
We conclude the paper by comparing the performance of QUBO solvers [31] (including Ising solvers, see [34]) against Relaxable Cardinality Constraints (RC2222) [26] — CDCL-based state-of-the-art Max SAT solver. [^3] RC2222 won the weighted and unweighted categories of the Max SAT Evaluation competition in $2018$ [37] and $2019$ [38]. Specifically, we focus on the accuracy, as measured by the number of violated clauses. The QUBO (and Ising) solvers use the gadget-based [27] conversion methods outlined in Section II (and Appendix B), whereas RC2222 solves the original $3$-SAT problem.
In Section IV.1, we used a set of four publicly available benchmark instances, while in Section IV.2, we used randomly generated instances (for details, see [35]). All of the tested instances in both sections are in the computationally hard regime, well beyond the difficulty phase transition point at clause density $\rho>1$, where $\rho=M/N$ with $M$ denoting the number of clauses and $N$ denoting the number of literals, see [32, 33].
For the QUBO solvers, we evaluated all $27$ heuristics listed in [31] and identified the “best” performing one, in this case the genetic algorithm with local search heuristics. Here, “best” is defined as the heuristic that produces the smallest median value. In instances where several heuristics yield identical median values, the one with the smallest minimum value is selected. For the Ising solver, we implemented ballistic simulated bifurcation (bSB) dynamics, utilizing the same parameter settings described in [34], but extending the number of time steps $\Delta_{t}$ to $20$ and maintaining a fixed $5000$ iterations. For performance comparison, we implemented the same selection strategy as outlined above. For RC2222, default parameters were employed. All results were generated in Ubuntu $20.04.6$ LTS on an Intel Core i7777-$1185$G7777 $3.0$GHz processor with $16$GB of RAM.
IV.1 Benchmark Comparison
We tested four sets of benchmark $3$-SAT instances from the Satisfiability Library [36] — uf50-218, AIM, PRET, and DUBIOS.
  • uf50-218 — we randomly selected $200$ instances out of $1000$ satisfiable ones in total. All of the tested instances are of size $N=50$ variables and $M=218$ clauses, corresponding to $\rho=4.4$.
  • AIM — we tested all $16$ satisfiable instances of size $N=50$ variables, where the number of clauses varies from $M=80$ to $300$; corresponding to clause density from $\rho=1.6$ to $6$.
  • PRET — we tested all $4$ unsatisfiable instances of size $N=60$ variables, $M=160$ clauses, and $\rho=2.7$.
  • DUBIOS — we tested $7$ unsatisfiable instances with varying sizes, from $N=60$ to $78$ variables and from $M=160$ to $208$ clauses (for all instances $\rho=2.7$).
The comparison of the performance of QUBO, Ising, and RC2222 solvers across the four sets of instances is shown in Fig. 1.
Figure 1: Benchmark. On the left side is a summary of the statistics of the performed numerical simulations. The highlighted values represent the best solutions for each of the four sets. The Q1, Q3, and Median values were calculated using the percentile function from NumPy (in Python). On the right side are the boxplots, corresponding to the instance with largest $N$ from each set. In case two instances from the same set have equal $N$, we selected the one with higher clause density $\rho$.
Evidently, QUBO matches the accuracy of RC2222 for almost all of the tested instances. For the Ising solver, however, we observe that it predominantly produces trivial solutions, such as all unnegated or negated variables, even after parameter tuning. Nevertheless, in certain instances from the PRET and DUBIOS sets, Ising achieves accuracy comparable to that of QUBO and RC2222. However, this only occurs because, in these specific cases, the trivial solutions coincide with the actual solutions.
IV.2 Clause Density
This Section provides additional insight into how the clause density $\rho$ affects the performance of QUBO solvers and RC2222. $\rho$ is well-recognized as an essential parameter determining the difficulty of $3$-SAT problems [32, 33, 39].
The simulations are conducted on randomly-generated instances of sizes $N=30,50,60,70$ variables. Here, each clause is constructed from exactly three distinct variables, picked uniformly at random, where each variable has $1/2$ probability of being negated. The clause density for each $N$ varies from $\rho=0.5$ to $6$ with a step size of $0.5$. For each $N$, we generate $10$ instances for every value of $\rho$. The instances are chosen so that we can finish the numerical simulations within a reasonable time while still being representative and covering both the easy and hard regimes of the problem. Indeed, as we can observe from Fig. 2, as $\rho$ increases in the given range, the problem exhibits a typical easy-to-hard pattern for all tested solvers, affecting their performance.
Figure 2: The number of violated clauses as a function of the clause density for four different instance sizes $N$. Each datapoint corresponds to 10 randomly-generated instances with 3 trials per instance.
In alignment with our previous results, we observe that, even for the hard instances (high clause density $\rho$), QUBO consistently matches the accuracy of RC2 across all sizes. What makes the results particularly appealing is the fact that popular alternative approaches, e.g. quantum approximate optimization algorithm (QAOA) [40] suffer from so-called reachability deficit (see, e.g., [41]) when dealing with $3$-SAT problems with high clause density $\rho$. Typically, this is attributed to the amount of frustration in the Ising variables in the cost Hamiltonian that encodes the problem. Consequently, QAOA has significant hardware overhead (especially in the hard region of the problem where the easy-to-hard pattern emerges), making it unsuitable for current NISQ devices. Despite continued attempts [42, 43, 44] to improve the capacity of QAOA to handle hard 3-SAT instances on near-term devices, its issues remain; further motivating the use of available QUBO solvers in the current hardware landscape.
The Ising solver, on the other hand, similar to subsection IV.1, cannot solve the randomly generated instances and produces mostly trivial solutions, i.e. all variables are either unnegated or negated. Unlike the results shown in Fig. 1, however, here the trivial solutions do not correspond to the actual solutions, and so the number of violated clauses is strictly bigger than zero even for $\rho=0.5$. Nonetheless, its trivial solutions do, in fact, satisfy some of the clauses, thus reducing the overall number of the violated clauses.
V Conclusions
This work paves the way towards establishing QUBO solvers as an alternative platform to CDCL-based solvers for tackling $3$-SAT problems, and it is the first performance-oriented comparison between them. To that end, we first demonstrated a $2$-step technique (using the $(7,10)$-gadget [27]) for mapping $3$-SAT problems of arbitrary size to the QUBO form via an intermediate step of converting them to their corresponding Max $2$-SAT formulations, and then deriving the corresponding QUBO cost function. The conversion is appealing as it allows us to retrieve the solution of the original $3$-SAT problem from the QUBO solution. This is necessary to enable the use of current QUBO (and Ising) solvers for $3$-SAT problems, see Appendix A and B.
We showed rigorously that the exact number of both satisfied and violated clauses of the original hard $3$-SAT problem can be computed explicitly from the converted Max $2$-SAT problem. Through numerical simulations, we demonstrated that available QUBO solvers can achieve state-of-the-art accuracy (matching that of RC2222) on several benchmark instances, even for the computationally hard cases with high clause density. Moreover, in the current NISQ era, the results have direct utility with regard to solving $3$-SAT problems on available quantum annealing hardware, and facilitate the development of advanced digital QUBO solvers and heuristics.
VI Data Availability
The entire source code and all of the raw data for the numerical analysis are openly available at [35].
Appendix A 3-SAT Hamiltonian does not have Ising quantization
Here we show that the 3-SAT Hamiltonian does not have an Ising quantization, necessitating the conversion approaches, outlined in Section II and Appendix B, to utilize QUBO and Ising solvers for tackling 3-SAT problems.
To ease the notation, for this Section we omit the upper index of the clause and let $S:=\left\{\sigma_{i}\right\}_{i=1}^{N}$ denote the Ising-spin literals of a 3-SAT problem with $M$ clauses $\left\{C_{\mu}\right\}_{\mu=1}^{M}$. That is for the following discussion we have $\sigma_{i}\in\left\{\pm 1\right\}$ and Eq. 2 of [45] reads:
$\mathcal{H}=\frac{M}{8}-\sum_{i=1}^{N}H_{i}\sigma_{i}-\sum_{i<j}T_{ij}\sigma_{i}\sigma_{j}-\sum_{i<j<k}J_{ijk}\sigma_{i}\sigma_{j}\sigma_{k},$ (9)
where $c_{\mu,i}=1$ if $\sigma_{i}$ is in clause $C_{\mu}$ and $c_{\mu,i}=-1$ otherwise. $H_{i}=\frac{1}{8}\sum_{\mu}c_{\mu,i},T_{ij}=-\frac{1}{8}\sum_{\mu}c_{\mu,i}\cdot c_{\mu,j},J_{ijk}=\frac{1}{8}\sum_{\mu}c_{\mu,i}\cdot c_{\mu,j}\cdot c_{\mu,k}$
Theorem A.1.
Quadratic penalty formulation of Eq. (9) is not possible for Ising-spin literals $\sigma_{i}\in\left\{\pm 1\right\}$.
Proof.
We begin by labeling the components of Eq. (9) as follows:
$\mathcal{H}=\frac{M}{8}-\underbrace{\sum_{i=1}^{N}H_{i}\sigma_{i}}_{\mathcal{L}}-\underbrace{\sum_{i<j}T_{ij}\sigma_{i}\sigma_{j}}_{\mathcal{Q}}-\underbrace{\sum_{i<j<k}J_{ijk}\sigma_{i}\sigma_{j}\sigma_{k}}_{\mathcal{C}}.$
The first three terms of $\mathcal{H}$ are at most quadratic, therefore it suffices to show that the cubic term $\mathcal{C}$ does not admit a quadratization over ±$\pm 1$ Ising spins. Suppose by contrary that $\mathcal{H}$ (equivalently $\mathcal{C}$) can be expressed as a quadratic polynomial over ±$\pm 1$ Ising spins with, at most, a quadratic penalty term. That is, for each $j,k=1,\ldots,N$, we can define an ancillary variable $y_{jk}\in\left\{-1,1\right\}$, a constant $\mu_{k}\in\mathbb{R}_{\geq 0}$, and a function $p_{k}:S\times S\times\left\{\pm 1\right\}\rightarrow\mathbb{R}$ such that:
  1. P.1
    $p_{k}$ is non-negative polynomial with $\operatorname{deg}(p_{k})\leq 2$
  2. P.2
    $p_{k}$ vanishes only when $\sigma_{j}\sigma_{k}-y_{jk}=0$ for each $k=1,\ldots,N$.
Eq. (9) thus becomes:
$\mathcal{H}=\frac{M}{8}$ $-\mathcal{L}-\mathcal{Q}$ (10)
$-\left(\sum_{i<j<k}J_{ijk}\sigma_{i}y_{jk}+\underbrace{\sum_{k=1}^{N}\mu_{k}p_{k}\left(\sigma_{j},\sigma_{k},y_{jk}\right)}_{\operatorname{Penalty}}\right).$
By P.1, $p_{k}$ is either linear or quadratic. The only linear polynomial (over the ancillary variable) that satisfies P.2 is $\sigma_{j}\sigma_{k}=y_{jk}$, which violates the non-negativity condition in P.1 when $\left(\sigma_{j},\sigma_{k},y_{jk}\right)=\left(1,-1,1\right)$. Therefore $p_{k}$ must be quadratic over all variables and there exists constants $A,B,C,D,E,F,G,H,I,J\in\mathbb{R}$ such that $p_{k}$ can be expanded into the following form:
$p_{k}$ $=B\sigma_{j}+C\sigma_{k}+Dy_{jk}$ (11)
$\quad+A\sigma_{j}\sigma_{k}+E\sigma_{j}y_{jk}+F\sigma_{k}y_{jk}$
$\quad+Gy_{jk}^{2}+H\sigma_{j}^{2}+I\sigma_{k}^{2}+J.$
In Table 3, we summarize the possible combinations of $\left(\sigma_{j},\sigma_{k},y_{jk}\right)$ and the corresponding coefficients of $f$ such that both P.1 and P.2 are satisfied.
$\sigma_{j}$ $\sigma_{k}$ $y_{jk}$ Constants when $p_{k}=0$:
$1$ $1$ $1$ $J=-A-B-C-D-E-F-G-H-I$
$-1$ $1$ $-1$ $J=A+B-C+D+E+F-H-H-I$
$1$ $-1$ $1$ $J=A-B+C+D+E-F-G-H-I$
$-1$ $-1$ $-1$ $J=-A+B+C-D-E+F-G-H-I$
Table 3: Choices of $\left(\sigma_{j},\sigma_{k},y_{jk}\right)$ and coefficients of $p_{k}$, such that both P.1 and P.2 are satisfied.
By P.2, we must have that all four rows of the right-most column of Table 3 are satisfied simultaneously. Hence $A=B=C=D=E=F=0$, and Eq. (11) reduces to:
$p_{k}=Gy_{jk}^{2}+H\sigma_{j}^{2}+I\sigma_{k}^{2}+J;\quad J=-G-H-I,$ (12)
which vanishes for all $\left(\sigma_{j},\sigma_{k},y_{jk}\right)\in S\times S\times\left\{\pm 1\right\}$ and P.2 is violated.
To see this, consider arbitrary Ising spin configuration $\left(\sigma_{j},\sigma_{k},y_{jk}\right)\in S\times S\times\left\{\pm 1\right\}$, Eq. (12) then becomes:
$p_{k}\left(\sigma_{j},\sigma_{k},y_{jk}\right)$ $=G\cdot 1+H\cdot 1+I\cdot 1+J$
$=G+H+I-G-H-I=0.$
This implies any quadratic penalty reduces to Eq. (12) necessarily violates P.2. We therefore conclude that a quadratic penalty $p_{k}$ that satisfies both P.1 and P.2 cannot exist, contradicting our claim that $\mathcal{H}$ can be expressed at most quadratically with such a penalty function and the proof is complete. ∎
By Theorem A.1, it follows immediately that a quadratic Hamiltonian formulation of the 3-SAT problem (e.g. as suggested by Eq. 4 of [46]) is not possible in the Ising spin variables $\sigma_{i}^{z}\in\left\{\pm 1\right\}$. The referenced equation is included below for completeness with $N$ replacing $k$ as the end point of the summation, denoting the number of variables, to avoid confusion:
$\mathcal{H}_{\text{clause}}^{(2)}$ $=\underbrace{J\cdot\sum_{j<i}^{N}c(i)c(j)\sigma_{i}^{z}\sigma_{j}^{z}}_{\text{Quadratic term}}+\underbrace{h\cdot\sum_{i=1}^{N}c(i)\sigma_{i}^{z}}_{\text{Linear term}}$
$+\underbrace{J^{a}\sum_{i=1}^{N}\sum_{j=1}^{N}c(i)\sigma_{i}^{z}\sigma_{j,a}^{z}}_{\text{Quad. coupled with aux from Cubic}}+\underbrace{\sum_{i=1}^{N}h_{i}^{a}\sigma_{i,a}^{z}}_{\text{aux. from Cubic}}$
Appendix B Max $2$-SAT to Ising
Similar to section II.2, we must construct a problem Hamiltonian $H_{\ddagger}$ [30] in order to utilize Ising solvers for Max $2$-SAT problems. Consider a Max $2$-SAT problem $\ddagger:=\bigwedge_{k=1}^{M^{\prime}}\Omega^{\prime}_{k}$ with $M^{\prime}$ clauses $\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M^{\prime}}$and $N^{\prime}$ literals $\left\{\sigma^{\prime}_{i}\right\}_{i=1}^{N^{\prime}}$. We now define a $H_{\ddagger}$ on $\ddagger$ over the search space $\mathbf{x}\in\left\{0,1\right\}^{N^{\prime}}$, for which we minimize over with the goal to minimize the total number of violated clauses. We first consider the negation binary variable $\mathbf{y}\in\left\{0,1\right\}^{N^{\prime}}$ of $\mathbf{x}$ as follows:
$y_{j}=1-x_{j},\quad\operatorname{for}j=1,\ldots,N^{\prime}.$
We then map binary variables $\{y_{j}\}_{j=1}^{N^{\prime}}$ to the ±$\pm 1$ eigenstates $\{\overline{\sigma}_{j}^{k}\}_{j=1}^{N^{\prime}}$ of the Pauli-$z$ operator:
$\overline{\sigma}_{j}^{k}\ket{y_{j}}=(-1)^{y_{j}}\ket{y_{j}}.$ (13)
We then define the set of intermediate variables $\{v_{j}^{k}\}$ over the $M^{\prime}$ clauses of $\ddagger$ as in section II.2.
Now we are ready to translate each $2$-literal clause $\Omega_{k}^{\prime}$ from the Max $2$-SAT problem $\ddagger$ into its corresponding $2$-local term $H_{\Omega_{k}^{\prime}}$
$H_{\Omega_{k}^{\prime}}=\frac{1-v_{j_{1}}^{k}\overline{\sigma}{j_{1}}^{k}}{2}\frac{1-v_{j_{2}}^{k}\overline{\sigma}_{j_{2}}^{k}}{2}$ (14)
The problem Hamiltonian $H_{\ddagger}$ can then be constructed in the following compact form
$H_{\ddagger}=\sum_{\Omega_{k}^{\prime}}H_{\Omega_{k}^{\prime}}$ (15)
where we simply sum over the energies of all $M^{\prime}$ clauses in the Max $2$-SAT problem. [^4] Interestingly, by virtue of the Hamiltonian (Eq. (15)) being the sum over all the energies of the $M^{\prime}$ clauses, its ground state corresponds to the solution of the Max $2$-SAT problem. Namely, the ground state could have positive energy, corresponding to the minimum number of violated clauses. This way, the converted Max $2$-SAT problem can be formulated in terms of the computational energy, where identical energy penalties of $1$ are assigned for every violated clause $H_{\Omega_{k}^{\prime}}$, and 00 for all satisfied ones. Expanding Eq. (15) then gives [^5] Notice the similarity with Eq. (4).:
$H_{\ddagger}=\frac{1}{4}\sum_{\Omega_{k}^{\prime}}1-v_{j_{1}}^{k}\overline{\sigma}^{k}_{j_{1}}-v_{j_{2}}^{k}\overline{\sigma}^{k}_{j_{2}}+v_{j_{1}}^{k}v_{j_{2}}^{k}\overline{\sigma}^{k}_{j_{1}}\overline{\sigma}^{k}_{j_{2}}$ (16)
where after rescaling by a factor of $4$ and dropping the first constant term we obtain, for each of the literals indexed by $j_{i}\in\left[1,\ldots,N^{\prime}\right]$, the entries of the external field $\left[h_{j_{i}}\right]\in\mathbb{R}^{N^{\prime}}$ and the coupling term $\left[J_{j_{1}j_{2}}\right]\in\mathbb{R}^{\left(N^{\prime}\right)\times\left(N^{\prime}\right)}$
$J_{j_{1}j_{2}}=\sum_{k=1}^{M^{\prime}}v_{j_{1}}^{k}v_{j_{2}}^{k}$ (17)
$h_{j_{i}}=-\sum_{k=1}^{M^{\prime}}v_{j_{i}}^{k}$ (18)
Notice that each of the entries in the Max $2$-SAT formulation, defined by the indices of the literals, is a sum over the set of $M^{\prime}$ clauses.
Appendix C Proof of Cor. III.3
Proof.
Suppose $\ddagger$ is solved when the ancillary variables satisfy $d^{k}=1$ for all $k=1,\ldots,M$. Let $\mathfrak{S}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses of $\ddagger$, and let $\mathcal{V},\mathcal{S}\in\mathbb{N}_{\geq 0}$ denote the number of violated and satisfied clauses of $\dagger$ respectively. We have the following linear Diophantine equation of three variables:
$\mathfrak{S}$ $=6\cdot\mathcal{S}_{1,2}+7\cdot\mathcal{S}_{1,3}+4\cdot\mathcal{V}$ (19)
$\mathcal{S}$ $=\mathcal{S}_{1,2}+\mathcal{S}_{1,3},$
where $\mathcal{S}_{1,2},\mathcal{S}_{1,3}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses under with ancillary variable $1$ in Case $2$ and Case $3$, respectively, in Table 2. Since $\operatorname{gcd}\left(4,6,7\right)=1$ and $\mathfrak{S}_{1}>1$, Eq. (19) has integral solutions obtained by applying Euclidean Algorithm twice on the following two linear Diophantine equations of two variables sequentially:
$\mathcal{W}$ $:=6\cdot\mathcal{S}_{1,2}+7\cdot\mathcal{S}_{1,3}$ (20)
$\mathfrak{S}$ $=\mathcal{W}+4\cdot\mathcal{V},$ (21)
and the proof is complete.
Appendix D Proof of Theorem III.4
Proof.
Let $I_{a}:=\left\{k\middle|d^{k}=a\right\}\subset\left\{1,\ldots,M\right\}$, then $I_{0}\sqcup I_{1}=\left\{1,\ldots,M\right\}$. Consider the partition of the set of converted clauses $\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M}$ of $\ddagger$ by the sets of indices $I_{0},I_{1}$: Let $\Omega^{\prime a}:=\left\{\Omega^{\prime}_{k}\right\}_{k\in I_{a}}$ for $a=0,1$, then $\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M}=\Omega^{\prime 0}\sqcup\Omega^{\prime 1}$.
Suppose $\ddagger$ is solved and let $\mathfrak{S}_{a}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses of $\Omega^{\prime a}\subset\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M}$ for $a=0,1$ respectively, and let $\mathcal{V}_{a},\mathcal{S}_{a}\in\mathbb{N}_{\geq 0}$ denote the corresponding number of violated and satisfied clauses of $\dagger$ for $a=0,1$ respectively.
In the case when the ancillary variables are $1$, i.e. in $\Omega^{\prime 1}\subset\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M}$, let $\mathcal{S}_{1,2},\mathcal{S}_{1,3}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses with ancillary variable $d^{k}=1$ under Case $2$ and Case $3$, respectively, in Table 2. Then $\mathcal{V}_{1},\mathcal{S}_{1}\in\mathbb{N}_{\geq 0}$ are integer solutions of the linear Diophantine equation:
$\mathfrak{S}_{1}$ $=6\cdot\mathcal{S}_{1,2}+7\cdot\mathcal{S}_{1,3}+4\cdot\mathcal{V}_{1}$
$\mathcal{S}_{1}$ $=\mathcal{S}_{1,2}+\mathcal{S}_{1,3},$
which we can compute by Corollary III.3.
Similarly for the subset of clauses where the ancillaries are 00, i.e. in $\Omega^{\prime 0}\subset\left\{\Omega^{\prime}_{k}\right\}_{k=1}^{M}$, $\mathcal{V}_{0},\mathcal{S}_{0}\in\mathbb{N}_{\geq 0}$ are integer solutions of the linear Diophantine equation:
$\mathfrak{S}_{0}=6\cdot\mathcal{V}_{0}+7\cdot\mathcal{S}_{0},$
Eq. (7) of Theorem III.2 can be written in terms of Case $2$ and Case $3$ of Table 2: Let $\mathcal{S}_{0,2},\mathcal{S}_{0,3}\in\mathbb{N}_{\geq 0}$ denote the total number of satisfied clauses with ancillary variable $d^{k}=0$ under Case $2$ and Case $3$, respectively, in Table 2, then by construction $\mathcal{S}_{0}=\mathcal{S}_{0,2}+\mathcal{S}_{0,3}$ and:
$\mathfrak{S}_{0}$ $=6\cdot\mathcal{V}_{0}+7\cdot\mathcal{S}_{0}$
$=6\cdot\mathcal{V}_{0}+7\cdot\mathcal{S}_{0,2}+7\cdot\mathcal{S}_{0,3}.$
which can be determined explicitly by Theorem III.2.
The total number of violated and satisfied clauses of $\dagger$ is thus given by $\mathcal{V}=\mathcal{V}_{0}+\mathcal{V}_{1}$ and $\mathcal{S}=\mathcal{S}_{0}+\mathcal{S}_{1}$, respectively, as desired.
To see that the sets are disjoint and, therefore, none of the pieces are being double counted, we enumerate them in Table 4.
$d=0$, Thm III.2 $d=1$, Cor III.3
Case 1 $6\cdot\underbrace{S_{0,1}}_{\mathcal{V}_{0}}$ $4\cdot\underbrace{S_{1,1}}_{\mathcal{V}_{1}}$
Case 2 $7\cdot\mathcal{S}_{0,2}$ $6\cdot\mathcal{S}_{1,2}$
Case 3 $7\cdot\mathcal{S}_{0,3}$ $7\cdot\mathcal{S}_{1,3}$
$+=\mathfrak{S}_{0}$ $+=\mathfrak{S}_{1}$
Table 4: Case by case satisfied clauses, notice the $d^{k}=0,1$ case are disjoint with the total number of satisfied clauses of $\dagger$ in each part being $\mathcal{S}_{0}=\mathcal{S}_{0,2}+\mathcal{S}_{0,3},\mathcal{S}_{1}=\mathcal{S}_{1,2}+\mathcal{S}_{1,3}$.